Today the sum of the ages of Susie and madly is 58.
In six years,four times Susie's age plus 10 will equal madly,s age at that time.
How old is each now.

Answer by ewatrrr(24785)

 
Hi
x is madley's age
x + 6 = 4((58-x) + 6)+10
x + 6 = 4(64 - x) + 10
5x = 260
x = 52, madley's age
Susie 6 years 58-52